Arcelor Mittal Dofasco can buy iron ore pellets from Minnesota at US$130 per short ton (2000 pounds) or from Labrador at C$175 per metric tonne (1000 kg). Which source is more expensive and by how much in C$ per metric tonne? (1 pound = 0.4536 kg, C$1.1400 = US$1.00) Weight and currency calculations should be to 4 decimal places; final answer rounded to the nearest cent.
Answer by Theo(13342)

exchange rate for C$ to US$ is 1.14 C$ = 1 US$.
price from minnesota is 130 US$ per short ton (2000 pounds).
price from labrador is 175 C$ per metric tonne (1000 kg).

1 pound = .4536 kg.
therefore 2000 pounds = .4536 * 2000 = 907.2 kg
1 US$ = 1.14 C$.
therefore 130 US$ = 148.2 C$.

price from minnesota is 130 US$ per 2000 pounds.
based on the US$ to C$ conversion factor and pound to kg conversion factor, ......
price from minnesota is 148.2 C$ per 907.2 kg.

148.2 C$ / 907.2 kg = x C$ / 1000 kg.
solve for x to get x = 148.2 / 907.2 * 1000 = 163.3597884 C$ per 1000 kg.

1 metric tonne is 1000 kg.
price from labrador is 175 C$ per metric tonne.
price from minnesota is 163.3597884 C$ per metric tonne.
price from labrador is 175 - 163.3597884 = 11.64021164 C$ per metric tonne more expensive = 11.6402 C$ per metric tonne more expensive (rounded to 4 decimal places).

that's what i get.